OpenStudy (anonymous):

how much would Adam have to deposit in a savings account to earn $100 in intrest over 4 years at 5% simple annual intrest rate?

OpenStudy (anonymous):

\[SI = \frac{ P \times R \times T }{ 1000 }\]

OpenStudy (anonymous):

SI = Simple Interest P=Principal (the money Adam would have to deposit in his savings account) R=Rate T=Time

OpenStudy (anonymous):

^^^^

OpenStudy (anonymous):

i need help!!!!!

OpenStudy (anonymous):

$500